# Scaling Down Deep Learning with MNIST-1D

### v4

@@ 1 Introduction
Meanwhile, MNIST has served as the initial proving ground for a large number of deep learning innovations including dropout, Adam, convolutional networks, generative adversarial networks, and variational autoencoders (Srivastava et al., 2014; Kingma and Ba, 2014; LeCun et al., 1989; Goodfellow et al., 2014; Kingma and Welling, 2013). Once a proof of concept was established in small-scale experiments, researchers were able to justify the time and resources needed for larger and more impactful applications.
@@ 2 Related work
There are a number of small-scale datasets that are commonly used to investigate science of deep learning questions. We have already alluded to MNIST (LeCun et al., 1998), CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009). The CIFAR datasets consist of $32\times 32$ colored natural images and have sample sizes similar to MNIST. They are better at discriminating between MLP and CNN architectures and also between different types of CNNs: for example, vanilla CNNs versus ResNets (He et al., 2015). The FashionMNIST dataset (Xiao et al., 2017) has the same size as MNIST and is somewhat more difficult. It aims to rectify some of the most serious problems with MNIST: in particular, that MNIST is too easy and thus all neural network models attain roughly the same test accuracy. None of these datasets is substantially smaller than MNIST and this hampers their use in fast-paced exploratory research or compute-heavy applications such as metalearning.
@@
To the best of our knowledge, the MNIST-1D dataset is unique in that it is over two orders of magnitude smaller than MNIST but can be used just as effectively — and in a number of important cases, more effectively — for studying fundamental deep learning questions. This may be why MNIST-1D was used as a teaching tool in the recent Understanding Deep Learning textbook (Prince, 2023).
@@ 3 The MNIST-1D dataset
Implementation. Our goal was to make the code as simple, modular, and extensible as possible. The code for generating the dataset occupies two Python files and fits in a total of 150 lines. The get_dataset method has a simple API for changing dataset features such as maximum digit translation, correlated noise scale, shear scale, final sequence length, and more (Table 2). The following code snippet shows the relative ease with which one can install the mnist1d package, choose a custom number of samples, and generate a dataset:
!pip install mnist1d  # get from pypi
from mnist1d.data import (make_dataset,
                      get_dataset_args)
default_args = get_dataset_args()
default_args.num_samples = 10000
data = make_dataset(default_args)
x,y,t = data["x"], data["y"], data["t"]
x.shape, y.shape, t.shape
>>> ((8000, 40), (8000,), (40,))
@@ 4 Classification
We used PyTorch to implement and train simple logistic, MLP, CNN, and GRU models. All models used the Adam optimizer and early stopping for model selection and evaluation. We obtained 32% accuracy with logistic regression, 68% using an MLP, 91% using a GRU, and 94% using a CNN (Table 1). Even though the test performance was markedly different between MLP, GRU, and CNN, all of them easily achieved 100% accuracy on the training set (Figure 2). While for the MLP this is a manifestation of overfitting, for the CNN this is an example of benign overfitting.
@@
We also trained the same models on a version of the dataset which was permuted along the spatial dimension. This ‘shuffled’ version measured each of the models’ performances in the absence of local spatial structure (Zhang et al., 2016; Li et al., 2018). The test accuracy of CNNs and GRUs decreased by about 35 percentage points after shuffling whereas the MLP and logistic models performed about the same (Table 1, Figure 2). This makes sense, as the former two models have spatial and temporal locality priors whereas the latter two do not.
@@ 5.1 Lottery tickets and spatial inductive biases
The original lottery ticket paper (Frankle and Carbin, 2019), as well as some of the follow-up studies (Morcos et al., 2019), required a large number of GPUs and multiple days of runtime. By contrast, all the experiments we presented here took around thirty minutes to complete on a single GPU.
@@ 5.2 Deep double descent
Figure 5: Deep double descent visible in the relationship between test error of an MLP classifier trained on MNIST-1D (+15% label noise) and the size of the hidden layer. The dashed line shows the location of the interpolation threshold. Adapted with permission from Prince (2023, Section 8.4). Runtime: $\sim$30 minutes. [CODE]
@@
Despite the above intuition, many aspects of double descent, such as what factors affect its width and location, are not well understood. We argue that MNIST-1D is well suited for exploring these questions. We observed double descent when training a MLP classifier on MNIST-1D, varying the size of the hidden layer. We observed test error peaking at the interpolation threshold (training error reaching zero), at around 50 neurons in the hidden layer (Figure 5). Further increasing the model size led to the test error dropping again. This experiment was adapted, with permission, from the Understanding Deep Learning textbook (Prince, 2023, Section 8.4) which uses the MNIST-1D dataset as a teaching aid. The runtime of this experiment was around 30 minutes on a single GPU.
@@ 5.5 Self-supervised learning
Figure 8: SimCLR-style (Chen et al., 2020) learning on MNIST-1D. (a) $t$-SNE embedding of the output representation after training ($n=5000$). (b) Linear classification accuracy after each layer. Layer 0 stands for the input (pixel space). Accuracy always peaks in the middle, in agreement with Bordes et al. (2023). Runtime: $\sim$5 minutes. [CODE]
@@ Acknowledgements
This work was partially funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via Germany’s Excellence Strategy (Excellence cluster 2064 “Machine Learning — New Perspectives for Science”, EXC 390727645; Excellence cluster 2181 “STRUCTURES”, EXC 390900948), the German Ministry of Science and Education (BMBF) via the Tübingen AI Center (01IS18039A) and the Gemeinnützige Hertie-Stiftung.

### v5

@@ 1 Introduction
Meanwhile, MNIST has served as the initial proving ground for a large number of deep learning innovations including dropout, Adam, convolutional networks, generative adversarial networks, and variational autoencoders (Srivastava et al., 2014; Kingma and Ba, 2014; LeCun et al., 1989; Goodfellow et al., 2014; Kingma and Welling, 2014). Once a proof of concept was established in small-scale experiments, researchers were able to justify the time and resources needed for larger and more impactful applications.
@@ 2 Related work
There are a number of small-scale datasets that are commonly used to investigate science of deep learning questions. We have already alluded to MNIST (LeCun et al., 1998), CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009). The CIFAR datasets consist of $32\times 32$ colored natural images and have sample sizes similar to MNIST. They are better at discriminating between MLP and CNN architectures and also between different types of CNNs: for example, vanilla CNNs versus ResNets (He et al., 2016). The FashionMNIST dataset (Xiao et al., 2017) has the same size as MNIST and is somewhat more difficult. It aims to rectify some of the most serious problems with MNIST: in particular, that MNIST is too easy and thus all neural network models attain roughly the same test accuracy. None of these datasets is substantially smaller than MNIST and this hampers their use in fast-paced exploratory research or compute-heavy applications such as metalearning.
@@
To the best of our knowledge, the MNIST-1D dataset is unique in that it is over two orders of magnitude smaller than MNIST but can be used just as effectively — and in a number of important cases, more effectively — for studying fundamental deep learning questions. This may be why MNIST-1D was used as a teaching tool in the recent Understanding Deep Learning textbook (Prince, 2023) [^1] The initial preprint of this manuscript was released on arXiv in November 2020..
@@ 3 The MNIST-1D dataset
Implementation. Our goal was to make the code as simple, modular, and extensible as possible. The code for generating the dataset occupies two Python files and fits in a total of 150 lines. The get_dataset method has a simple API for changing dataset features such as maximum digit translation, correlated noise scale, shear scale, final sequence length, and more (Table 2). The following code snippet shows how to install the mnist1d package, choose a custom number of samples, and generate a dataset:
# install the package from PyPI
# pip install mnist1d
from mnist1d.data import make_dataset
from mnist1d.data import get_dataset_args
args = get_dataset_args() # default params
args.num_samples = 10_000
data = make_dataset(args)
x, y = data["x"], data["y"]
The frozen dataset with $4000+1000$ samples can be found on GitHub as mnist1d_data.pkl.
@@ 4 Classification
We used PyTorch to implement and train simple logistic, MLP (fully-connected), CNN (with 1D convolutions), and GRU (gated recurrent unit) models. We used the Adam optimizer and early stopping for model selection and evaluation. We obtained 32% accuracy with logistic regression, 68% using an MLP, 91% using a GRU, and 94% using a CNN (Table 1). Even though the test performance was markedly different between MLP, GRU, and CNN, all of them easily achieved 100% accuracy on the training set (Figure 2). While for the MLP this is a manifestation of overfitting, for the CNN this is an example of benign overfitting.
@@
We also trained the same models on a version of the dataset which was permuted along the spatial dimension. This ‘shuffled’ version measured each of the models’ performances in the absence of local spatial structure (Zhang et al., 2017; Li et al., 2018). The test accuracy of CNNs and GRUs decreased by about 35 percentage points after shuffling whereas the MLP and logistic models performed about the same (Table 1, Figure 2). This makes sense, as the former two models have spatial and temporal locality priors whereas the latter two do not.
@@ 5.1 Lottery tickets and spatial inductive biases
The original lottery ticket paper (Frankle and Carbin, 2019), as well as some of the follow-up studies (Morcos et al., 2019), required a large number of GPUs and multiple days of runtime. By contrast, all the experiments we presented here took around $\sim$30 minutes to complete on a single GPU.
@@ 5.2 Deep double descent
Figure 5: Deep double descent in MNIST-1D classification. Here the test set had 12 00012\,000 samples. (a) MLP classifier with one hidden layer. (b) MLP classifier; 15% label noise. (c) CNN classifier with three convolutional layers; 15% label noise. Adapted with permission from Prince (2023, Section 8.4). CPU runtime: $\sim$60 minutes. [CODE]
@@
Despite the above intuition, many aspects of double descent, such as what factors affect its width and location, are not well understood. We argue that MNIST-1D is well suited for exploring these questions. We observed double descent when training a MLP classifier on MNIST-1D, varying the size of the single hidden layer. In the presence of 15% label noise, the test error peaked at the interpolation threshold (training error reaching zero), at around 50 neurons in the hidden layer (Figure 5b). Further increasing the model size led to the test error dropping again. Without label noise, the test error did not peak (Figure 5a). We observed qualitatively similar behavior using the CNN architecture (Figure 5c). The runtime of this experiment was $\sim$60 minutes on a CPU.
@@ 5.5 Self-supervised learning
Figure 8: SimCLR-style (Chen et al., 2020) learning on MNIST-1D. (a) $t$-SNE embedding of the output representation after training ($n=5000$). (b) Linear classification accuracy after each layer. Layer 0 stands for the input (pixel space). Accuracy always peaks in the middle (Bordes et al., 2023). CPU runtime: $\sim$5 minutes. [CODE]
@@ Acknowledgements
This work was partially funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via Germany’s Excellence Strategy (Excellence cluster 2064 “Machine Learning — New Perspectives for Science”, EXC 390727645; Excellence cluster 2181 “STRUCTURES”, EXC 390900948), the German Ministry of Science and Education (BMBF) via the Tübingen AI Center (01IS18039A), and the Gemeinnützige Hertie-Stiftung.